OpenStudy (anonymous):

@rizags @acxbox22 help me factor 3p+1=p^2+3p^3

OpenStudy (anonymous):

first move everything to one side of the equation

OpenStudy (anonymous):

ok so 3p+1-p^2-3p^3=0 right?

OpenStudy (anonymous):

i can't...

OpenStudy (anonymous):

they need to have the same exponents

OpenStudy (trojanpoem):

My bad. I saw them p^3 and 3p^3 Now multiply the whole equation by -1 and arrange them from higher exponent to lower exponent.

OpenStudy (anonymous):

why do i multiply it by a negative 1

OpenStudy (trojanpoem):

It looks better as 3p^3 + p^2 - p -1 = 0 well, never mind.

OpenStudy (trojanpoem):

now take -p^2 as a common factor from 3p^3 + p^2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

can you just write out all the steps to get the answer

OpenStudy (trojanpoem):

-p^2 ( -p -1) - p - 1 = 0 now take (-p -1 ) (-p -1) ( - p^2 + 1) = 0 * -1 (p+1)(p^2 - 1) = 0 now factorize (p^2 -1 ) (p+1)(p+1)(p-1) = 0 p = 1 , -1

OpenStudy (anonymous):

awesome thanks! can i ask some more questions on different problems?

OpenStudy (trojanpoem):

Sure. But did you understand the answer ?

OpenStudy (anonymous):

yes and no. i dont understand how you got the first part...

OpenStudy (trojanpoem):

which part exactly ?

OpenStudy (anonymous):

what im getting is (3p +1) (-p^2+1) making p= 1/3 and 1

OpenStudy (trojanpoem):

Well, assume we have x^3 + x^2 + x + 1 = 0 we want to factorize it. So our ambition to get two similar () ex: (y+x) + (y+x) = (y+x)(1+!1). x^3 + x^2 = -x - 1 x^3(x+1) = -x -1 now get -x -1 back to the other side x^3(x+1) + (x +1) = 0 the same as 2 +2 = 2(1+1) so (x+1)(x^3 +1 ) = 0

OpenStudy (anonymous):

but thats not the math problem , its 3p+1=p^2+3p^3

OpenStudy (trojanpoem):

I am bad at explaining. Check this : http://www.wikihow.com/Factor-a-Cubic-Polynomial

OpenStudy (trojanpoem):

I was just showing you example , and you have to apply what you understand on your equation even the first solution was without the (3)

OpenStudy (anonymous):

ok i understand your intentions but the confused me more can u at least solve this? m^2 +2mn +n^2-16p^2 i think its prime...?

OpenStudy (trojanpoem):

n , p , m ?

OpenStudy (trojanpoem):

m^2 + 2mn + n^2 = 16p^2 (m + n)(m+n) - 16p^2= 0 (4p - (m+n))(4p + (m+n))

OpenStudy (trojanpoem):

(m + n)(m+n) - 16p^2 <<< 16p^2 - (m+n)^2 ( difference between two squares)

OpenStudy (anonymous):

m^2 +2mn +n^2-16p^2 its subtracting 16...?

OpenStudy (trojanpoem):

I know , factorize m^2 + 2mn + n^2 without looking at -16p^2 or get it to the other side m^2 + 2mn + n^2 = 16p^2 same as zero the only difference is that you can't saw (m+n) = 16p^2

OpenStudy (trojanpoem):

saw -> say

OpenStudy (anonymous):

ok i really appreciate your help, but can you give me an anser so we can then talk about how we got it?

OpenStudy (trojanpoem):

p = m+n/2 p = -m-n/2

OpenStudy (anonymous):

ok, so how did that happen?

OpenStudy (trojanpoem):

m^2 +2mn +n^2-16p^2 = 0 m^2 + 2nm + n^2 <-- Like any other equation but with symbols instead of numbers